a truck is traveling on a straight line on ground level and is accelerating uniformly with an acceleration of magnitude (a). a rope is tied to a bucket of mass (m), the bucket tosses widely when the truck starte to accelerate but later on displaces at a fixed position at a fixed distance behind the truck and the rope hanging straight at a fixed angle; find the angle at which the rope will settle.express your answers in terms of m, g and a
1
Expert's answer
2020-02-10T09:34:24-0500
"T\sin{\theta}=ma"

"T\cos{\theta}=mg"

Thus,

"\theta=\arctan{\frac{a}{g}}"